OpenStudy (anonymous):

a swimmer runs over a horizontal diving board at a speed of 0.95 meters/second. 2.0 seconds after she dives, she hits the water. what was the height of the diving board from the waters surface and how far from the board was the swimmer when she hit the water?

OpenStudy (abmon98):

Horizontal diving board-0.95 m/s Vertical Velocity=0 m/s if the swimmer is released from rest under gravity. Vertical Displacement=1/2at^2 Horizontal Displacement=u*t